The following table gives the map coordinates and the shipping loads for a set of cities that we wish to connect through a cehow to solve b) X and Y
Answer #1

Solution:-

Part:- A

City

X

Y

Shipping Load

Weighted X

Weighted Y

A

4

10

5

20

50

B

5

8

11

55

88

C

4

9

15

60

135

D

7

6

5

35

30

E

6

9

16

96

144

F

3

3

9

27

27

G

3

5

4

12

20

Total

65

305

494

X bar = 305 / 65 = 4.69

Y bar = 494 / 65 = 7.6

Part B:-

We need to change the quantity of shipping load of City A from 5 to 15 (Triple)

City

X

Y

Shipping Load

Weighted X

Weighted Y

A

4

10

15

60

150

B

5

8

11

55

88

C

4

9

15

60

135

D

7

6

5

35

30

E

6

9

16

96

144

F

3

3

9

27

27

G

3

5

4

12

20

Total

75

345

594

X bar = 345 / 75 = 4.6

Y bar = 594 / 75 = 7.92
Answer #2

The following table gives the map coordinates and the shipping loads for a set of cities that we wish to connect through a central hub.  

CityMap Coordinates (x.y)Shipping Load
A(6,11)6
B(7,9)9
C(5,8)14
D(8,4)5
E(8,8)15
F(4,3)11

a) For the location of the proposed new central hub, the coordinates should be near:

x = ___ (round your response to two decimal places).

y = ___ (round your response to two decimal places).

b) If the shipment from city A triple, for the location of the proposed new central hub, the coordinate should be near:

x = ___ (round your response to two decimal places).

y = ___ (round your response to two decimal places).


source: Operations Management
answered by: Amber